OpenStudy (anonymous):

What are the x-intercepts of the graph of y=12x^2-5x-2? A) 1 and -1/6 B) -1 and 1/6 C) 2/3 and -1/4 D) -2/3 and 1/4

OpenStudy (anonymous):

the x-intercept is when y=0. so basically you will want to factor the right side and find it's two zeroes.

OpenStudy (anonymous):

so quadratic formula?

OpenStudy (anonymous):

i mean completing the square

OpenStudy (anonymous):

please help

OpenStudy (callisto):

quadratic formula will work.

OpenStudy (anonymous):

i completed the square... but im confused where to go from there

OpenStudy (callisto):

Can you show what you've got for completing square?

OpenStudy (anonymous):

I got down to this y+2= (x-5/2)^2

OpenStudy (anonymous):

how dow i find the x-ints.?

OpenStudy (callisto):

That doesn't look right. How did you get it?

OpenStudy (callisto):

To find x intercepts, you put y=0 and solve for x.

OpenStudy (anonymous):

y+2=1(12x^2-5x+25/4) y+2=(x-5/2)^2

OpenStudy (anonymous):

how do I find two x-ints? one I plug in 0 but then the other?

OpenStudy (callisto):

1. That's not the right way to do completing square. 2. You can find two x-ints by putting y=0 into the equation, since the degree of y is 2 and delta is not 0. You need to solve x.

OpenStudy (callisto):

To do completing square, you need to factorize the first two term and pick the coefficient of x^2 as the common factor, that is 12 in this case. \[y=12x^2-5x-2\]\[y=12(x^2-\frac{5}{12}x)-2\]Then, you need to add and subtract the same term inside the bracket. What would it be?

OpenStudy (anonymous):

so I did the quadratic formula instead and got (5plus or minus 11)/24

OpenStudy (callisto):

Looks good. So, the first x would be (5+11)/24, and the second x would be (5-11)/24

OpenStudy (anonymous):

isn't that answer what y equals?

OpenStudy (callisto):

No, that's what x equals to. Since you've put y=0 to solve x.

OpenStudy (anonymous):

wait.. isn't it a y equals equation? kinda confused

OpenStudy (callisto):

y=12x^2-5x-2 To find x-ints, put y=0. So, 0=12x^2-5x-2 Using quadratic formula, \[x= \frac{-(-5)\pm \sqrt{(-5)^2 - 4(12)(-2)}}{2(12)}\] It's the x which you're solving using the quadratic formula.

OpenStudy (anonymous):

so the x-ints would be -1/4 and 2/3 right

OpenStudy (callisto):

Yes.

OpenStudy (anonymous):

A plot is attached.